OpenStudy (anonymous):

What is the measure of the smaller angle formed by the hour and minute hands at 1:28? Please explain.

hero (hero):

NO that's not right either

OpenStudy (anonymous):

I am in 8th grade. is there any simple method

hero (hero):

Well, that is supposed to be the simple method

hero (hero):

let me see.. I think I have it

OpenStudy (anonymous):

can you explain Please

hero (hero):

i will

hero (hero):

The answer is 138 degrees

hero (hero):

Okay, basically, you find the angle for the hour hand. To find that use the following formula: 1/12 = x/360 you get x = 30 degrees Then you find the angle for the minute hand: 28/60 = x/360 you get x = 168 degrees Then you subtract 168 - 30 to get 138 degrees

OpenStudy (anonymous):

Sorry, I wasn't here for a while .Thank you very much for your help!